OpenStudy (anonymous):

Need help on the following question, Iodine-123, used in thyroid therapy, has a half-life of 13.27 hours. How many half-lives are required for a 160 mg sample of iodine-123 to decay to 5.0 mg? A) 1.0 B) 32 C) 5.0 D) 0.031

OpenStudy (anonymous):

t = 0 => 160 mg t = 13.27 hrs => 80 mg t = 13.27 * 2 hrs => 40 mg t = 13.27 * 3 hrs => 20 mg t = 13.27 * 4 hrs => 10 mg t = 13.27 * 5 hrs => 5 mg So after 66.35 hrs, i.e. after 5 halflives, there will be only 5 mg left.